OpenStudy (anonymous):

y = sinh-1(7x) how to difrentiate?

OpenStudy (nowhereman):

Use the formula for differentiating inverse functions.

OpenStudy (anonymous):

dy/dx=1/cosh(7x) is that correct?

OpenStudy (nowhereman):

No, the correct formula is: \[\frac{df^{-1}(y)}{dy} = \frac{1}{\frac{df}{dx}(f^{-1}(y))}\]